On the first day of school, the percentage of boys in a particular class is 60%. During the school year, six girls move away, and are replaced in the class by six boys; this makes the class roster 75% boys. Find the number of boys and girls in the class on the first day of school.
Answer by rwm(914)

b=boys
g=girls
c=b+g
b/c=.6
b+6/c=.75
24 boys and 16 girls to start
check
24/40=.6
30/40=.75